### v4

# Towards Holistic Evaluation of Large Audio-Language Models: A Comprehensive Survey

Abstract
With advancements in large audio-language models (LALMs), which enhance large language models (LLMs) with auditory capabilities, these models are expected to demonstrate universal proficiency across various auditory tasks. While numerous benchmarks have emerged to assess LALMs’ performance, they remain fragmented and lack a structured taxonomy. To bridge this gap, we conduct a comprehensive survey and propose a systematic taxonomy for LALM evaluations, categorizing them into four dimensions based on their objectives: (1) General Auditory Awareness and Processing, (2) Knowledge and Reasoning, (3) Dialogue-oriented Ability, and (4) Fairness, Safety, and Trustworthiness. We provide detailed overviews within each category and highlight challenges in this field, offering insights into promising future directions. To the best of our knowledge, this is the first survey specifically focused on the evaluations of LALMs, providing clear guidelines for the community. We will release the collection of the surveyed papers and actively maintain it to support ongoing advancements in the field.
Towards Holistic Evaluation of Large Audio-Language Models: A Comprehensive Survey
Chih-Kai Yang1, Neo S. Ho1, Hung-yi Lee1,2 1National Taiwan University, 2NTU Artificial Intelligence Center of Research Excellence (NTU AI-CoRE) chihkaiyang1124@gmail.com https://github.com/ckyang1124/LALM-Evaluation-Survey
1 Introduction
Recent advancements in large language models (LLMs) (Zhao et al., 2023; Grattafiori et al., 2024; Hurst et al., 2024) have expanded their impact beyond natural language processing (NLP) to multimodal domains (Yin et al., 2024; Team et al., 2024). Among these, large audio-language models (LALMs) (Lakhotia et al., 2021; Tang et al., 2024; Chu et al., 2024; Lu et al., 2024; Défossez et al., 2024; Fang et al., 2025) have attracted significant attention in the auditory-processing community. LALMs are multimodal LLMs that process auditory and/or textual input, such as speech, audio, and music, and generate textual and/or auditory output. They can be trained from scratch or fine-tuned from text LLM backbones with auditory modalities inserted. By integrating auditory modalities with language understanding, they show potential in auditory processing (Huang et al., 2024a), multimodal reasoning (Sakshi et al., 2025), and human-computer interaction (Lin et al., 2024a).
As LALMs evolve, expectations for their capabilities have expanded from basic tasks like speech recognition to more complex ones such as audio-grounded reasoning (Sakshi et al., 2025) and interactive dialogue (Lin et al., 2025a). Figure 1 illustrates this multifaceted nature, emphasizing the diverse input and output modalities involved and the wide range of abilities these models are expected to demonstrate. To evaluate these capabilities, a variety of benchmarks have been developed (Lin et al., 2025a; Yang et al., 2024c; Cheng et al., 2025).
However, the evaluation landscape remains fragmented and lacks systematic organization. Existing surveys (Wu et al., 2024a; Peng et al., 2024; Cui et al., 2024; Arora et al., 2025a) focus primarily on model architectures and training methodologies, with less emphasis on the equally important role of evaluation in assessing LALMs’ capabilities. This gap makes it challenging for researchers to find suitable benchmarks for their models or to pinpoint the field’s progress. Therefore, a structured overview of LALM evaluation frameworks is needed.
This paper presents a comprehensive survey of LALM evaluation frameworks and introduces a taxonomy categorizing evaluation dimensions. To the best of our knowledge, this is the first in-depth survey and taxonomy specifically focused on LALM evaluation. We organize the frameworks into four primary categories: General Auditory Awareness and Processing Section˜3, Knowledge and Reasoning Section˜4, Dialogue-oriented Ability Section˜5, and Fairness, Safety, and Trustworthiness Section˜6. We also highlight challenges in LALM evaluation section˜7, such as data contamination and insufficient consideration of human diversity, while suggesting promising future directions.
Overall, our contributions are threefold: (1) presenting the first comprehensive survey of LALM evaluations, (2) proposing a structured taxonomy for LALM evaluation that offers clear guidelines for researchers, and (3) identifying key challenges and future directions to improve evaluation coverage and robustness.
2 Taxonomy of Evaluation Frameworks for Large Audio-Language Models
As LALMs integrate multimodal understanding, they tackle tasks across speech, audio, and music. Despite numerous benchmarks for LALMs emerging, the evaluation landscape remains fragmented. To address this, we present the first structured taxonomy of LALM evaluations.
Figure 2 shows our taxonomy, with some works included. The full categorization of the surveyed works is in Appendix A. We organize the surveyed works into four categories by evaluation objectives:
General Auditory Awareness and Processing evaluates the auditory awareness and fundamental processing tasks, e.g., speech recognition and audio captioning.
Knowledge and Reasoning assesses LALMs’ knowledge acquisition and advanced reasoning skills, examining their intelligence.
Dialogue-oriented Ability focuses on natural conversational skills, including affective and contextual interaction, dialogue management, and instruction following.
Fairness, Safety and Trustworthiness examines bias, toxicity, and reliability for ethical, safe, and trustworthy deployment.
Each category is further divided into subcategories, as shown in Figure 2. Please note that, since existing benchmarks are inherently multi-dimensional, some are listed under multiple categories due to their multifaceted design. This typically happens in two cases: (1) when a benchmark comprises multiple tasks that independently assess different capabilities (e.g., VoiceBench includes tasks for both world knowledge and safety evaluation), and (2) when a single task requires the integration of several skills (e.g., certain MMAU tasks demand both expert knowledge and reasoning ability).
The following sections provide a detailed overview, highlighting the current progress, limitations, and future directions.
3 General Auditory Awareness and Processing
A distinctive strength of LALMs over cascaded systems (Huang et al., 2024c; Kuan et al., 2024b) is their inherent ability to directly interpret auditory signals, capturing crucial non-verbal cues such as speaker identity, emotion, and ambient context, without relying on separate components like speech recognition or emotion recognition systems connected to an LLM. This section reviews works evaluating both acoustic awareness and foundational auditory processing, emphasizing these core capabilities that set LALMs apart from LLMs.
3.1 Auditory Awareness
Benchmarks for auditory awareness examine how effectively LALMs realize acoustic cues like emotion, prosody, and environmental sounds. SALMon (Maimon et al., 2025) specifically evaluates sensitivity to acoustic inconsistencies (e.g., sudden speaker or emotional changes) and misalignments between acoustic signals and semantic content (e.g., conveying sad content with a cheerful tone). These evaluations reveal significant gaps between LALMs and human-level perception.
EmphAssess (Seyssel et al., 2024) measures LALMs’ awareness of prosodic emphasis by requiring speech-to-speech paraphrasing or translation that accurately preserves and transfers emphasis on specific parts of the input utterance. This evaluates LALMs’ ability to capture and maintain fine-grained prosodic features.
3.2 Auditory Processing
Building on auditory awareness, LALMs must also excel in fundamental auditory tasks, such as speech recognition, audio classification, and music analysis, to support advanced real-world applications. A list of commonly evaluated tasks and their corresponding datasets is provided in Appendix C for reference. Initially driven by representation learning models (Baevski et al., 2020; Hsu et al., 2021; Li et al., 2022), enriched datasets (Pratap et al., 2020; Piczak, 2015a; Hawthorne et al., 2019), and existing benchmarks (Yang et al., 2021; Turian et al., 2022; Yuan et al., 2023), recent works adapt these resources into instruction-oriented evaluation frameworks tailored for LALMs.
Dynamic-SUPERB (Huang et al., 2024a) initiated this direction, constructing 55 multiple-choice question-answering (QA) tasks spanning speech, audio, and music modalities. Subsequent efforts, such as AIR-Bench (Yang et al., 2024c) and AudioBench (Wang et al., 2025a), extend to open-ended QA formats. MuChoMusic (Weck et al., 2024) specifically emphasizes music-related tasks, while Dynamic-SUPERB Phase-2 (Huang et al., 2025a) significantly enlarges the benchmark to 180 tasks, forming the largest evaluation suite for LALMs’ general processing abilities to date.
Given the task diversity, various evaluation metrics are adopted depending on the task specificity, such as word error rate for speech recognition and BLEU score (Papineni et al., 2002) for translation. There is also an emerging trend that includes LLM-as-a-judge (Gu et al., 2024) for scalable, automatic evaluation of open-ended responses (Huang et al., 2025a; Yang et al., 2024c; Wang et al., 2025a).
Despite achieving promising results in certain areas, these benchmarks demonstrate that current LALMs still fall short of universally robust performance across auditory-processing tasks (Huang et al., 2025a), highlighting substantial room for improvement toward truly auditory foundation models.
4 Knowledge and Reasoning
Intelligent LALMs should demonstrate extensive knowledge and advanced reasoning to tackle complex real-world tasks. Current evaluations emphasize these abilities through three categories: Linguistic Knowledge, World Knowledge Assessment, and Reasoning. Each category targets distinct but complementary skills, collectively providing a comprehensive evaluation. These assessments reveal key challenges LALMs face in mastering knowledge and reasoning for advanced tasks.
4.1 Linguistic Knowledge
Linguistic knowledge refers to understanding and effectively using spoken language. Evaluating LALMs’ linguistic proficiency typically use likelihood-based benchmarks where models choose the more linguistically plausible option from paired speech samples. These tests cover lexical knowledge, syntax, and semantic coherence.
Representative works include the ZeroSpeech 2021 benchmark (Nguyen et al., 2020), which consists of multiple tracks for evaluating linguistic capabilities. The lexical-level assessment track, sWUGGY, tests models’ ability to distinguish between real words and phonotactically similar nonwords, while the syntactic sensitivity evaluation track, sBLIMP, focuses on differentiating grammatical from ungrammatical sentences. CSZS (Huang et al., 2024b) extends syntactic evaluation to multilingual and code-switched scenarios. Narrative and semantic coherence are evaluated by tasks like sStoryCloze and tStoryCloze (Hassid et al., 2023), where models are tasked with selecting semantically appropriate continuations to spoken stories.
4.2 World Knowledge Assessment
Real-world tasks often demand integrating external knowledge beyond basic auditory understanding. World knowledge assessment evaluates LALMs on two main aspects: (1) auditory expertise like music structure and medical sound diagnosis, and (2) general commonsense and factual knowledge.
Benchmarks that evaluate auditory expertise include MuChoMusic (Weck et al., 2024) and MMAU (Sakshi et al., 2025), which focus on musical understanding, such as melodic structure, harmony, instrument identification, and contextual music interpretation. Additionally, SAGI (Bu et al., 2024) assesses medical expertise, such as recognizing illnesses from audio cues like coughing.
Commonsense and factual knowledge evaluations often convert established text benchmarks into spoken form using text-to-speech (TTS). VoxEval (Cui et al., 2025) and VoiceBench serve as spoken counterparts to MMLU (Hendrycks et al., 2021) and MMLU-Pro (Wang et al., 2024), testing models across diverse factual domains like social science and humanities. Audiopedia (Penamakuri et al., 2025) uses knowledge graphs from Wikidata (Vrandečić and Krötzsch, 2014) to create audio-based, knowledge-intensive QA tasks that evaluate models’ knowledge of well-known entities, such as brands, mentioned in audio.
These benchmarks thoroughly assess LALMs’ knowledge acquisition, revealing challenges such as limited auditory expertise (Weck et al., 2024) and inconsistent performance across domains. Different LALMs excel in different domains, each with their own strengths, but their performance often noticeably declines outside their own specialized areas (Cui et al., 2025). Overall, there remains substantial room to improve LALMs’ auditory expertise and factual knowledge.
4.3 Reasoning
Reasoning over auditory inputs falls into two types. Content-based reasoning tests a model’s ability to understand spoken semantic content and answer questions. Acoustic-based reasoning requires utilizing acoustic features like speaker traits and environmental sounds beyond semantics. We provide an overview of these two evaluation paradigms.
4.3.1 Content-based Reasoning
Content-based reasoning assesses LALMs’ ability to reason over the semantic content of auditory queries. Current benchmarks for this capability typically transform NLP reasoning benchmarks into spoken questions via TTS and require LALMs to provide answers. For instance, VoxEval (Cui et al., 2025), URO-Bench (Yan et al., 2025), and ADU-Bench (Gao et al., 2024) convert NLP datasets like GSM8K (Cobbe et al., 2021) and MMLU (Hendrycks et al., 2021) into speech, evaluating LALMs’ mathematical reasoning based on spoken questions. During synthesis, various speaking styles (e.g., mispronunciation, disfluencies, and accents) may be introduced to test models’ robustness (Cui et al., 2025).
4.3.2 Acoustic-based Reasoning
Acoustic-based reasoning requires LALMs to infer from acoustic cues in auditory input, often involving reasoning across multiple auditory modalities or combining auditory understanding with cognitive skills such as compositional, temporal, logical, and multi-hop reasoning.
Cross-auditory Modality Reasoning demands joint reasoning over multiple auditory modalities, like speech and non-speech sounds. Wang et al. (2025d) propose an open-ended QA benchmark assessing co-reasoning on speech and environmental sounds, requiring reasoning over cues from distinct auditory sources to infer speakers’ activities. Their findings show that current LALMs frequently neglect non-speech cues, leading to failures.
Compositional and Temporal Reasoning involves comprehending structured acoustic events, their temporal relationships, and attribute binding. Benchmarks like CompA (Ghosh et al., 2024b) evaluate these abilities through specific tasks: CompA-order challenges models to identify correct event sequences or align audio temporal structures with textual descriptions, while CompA-attribute focuses on associating sound events with their sources and attributes. MMAU (Sakshi et al., 2025) assesses temporal reasoning via event counting and duration comparison.
Logical reasoning covers structured inference, including deductive and causal reasoning. Deductive reasoning can be tested by Audio Entailment (Deshmukh et al., 2025a), which evaluates whether a textual hypothesis logically follows from auditory input based on acoustic attributes like sound sources. MMAU (Sakshi et al., 2025) examines LALMs’ causal reasoning on cause-and-effect relationships of events.
Multi-hop reasoning is the ability to recall and integrate multiple information to answer complex queries, enabling models to connect stored knowledge without explicit reasoning steps (Yang et al., 2024d, e; Biran et al., 2024). SAKURA (Yang et al., 2025a) evaluates LALMs’ multi-hop reasoning by requiring integration of auditory attributes (e.g., speaker gender and emotion) with stored knowledge. Results show that LALMs struggle to combine auditory information with stored knowledge for reasoning, even when both types of information are extracted and known by the models.
5 Dialogue-oriented Ability
While foundational skills such as auditory awareness Section˜3.1, fundamental processing Section˜3.2, language proficiency Section˜4.1, advanced knowledge Section˜4.2, and reasoning Section˜4.3 are essential for LALMs, natural human-AI interactions additionally require affective and contextual interaction, fluent dialogue management, and precise instruction following. This category targets these integrative skills, focusing on naturalness and controllability, which we group as Conversational Ability and Instruction Following.
5.1 Conversational Ability
Effective conversational ability in LALMs relies on generating contextually appropriate responses and smoothly managing dialogues in real time. Current evaluations address this via two complementary frameworks: affective and contextual interaction, and full-duplex dialogue management.
5.1.1 Affective and Contextual Interaction
Evaluations of affective and contextual interaction typically adopt half-duplex settings, focusing on fully turn-by-turn conversations without speaker overlaps. These benchmarks emphasize LALMs’ ability to respond using both content and non-content cues such as emotional tone, speaking style, and speaker traits. StyleTalk (Lin et al., 2024a) presents models with a dialogue history and the user’s current speech segment, intentionally leaving the user’s intent underspecified when relying solely on the content. Consequently, models are required to leverage paralinguistic cues to respond appropriately. Subsequent works, such as SD-Eval (Ao et al., 2024) and VoxDialogue (Cheng et al., 2025), broaden the evaluation by incorporating more acoustic and contextual variables, including speaker age, accent, and environmental conditions. These benchmarks combine objective metrics (e.g., ROUGE-L (Lin, 2004), METEOR (Banerjee and Lavie, 2005)), LLM-based judgment (Gu et al., 2024), and human evaluation for comprehensive assessment.
5.1.2 Full-duplex Dialogue Management
Full-duplex evaluation examines LALMs in real-time, dynamic dialogues with complex behaviors like turn-taking (Duncan, 1972; Gravano and Hirschberg, 2011), backchanneling (Schegloff, 1982), and speaker interruptions and overlaps (Gravano and Hirschberg, 2012; Schegloff, 2000). These behaviors are detailed in Appendix F.
Representative works, such as Talking Turns (Arora et al., 2025b) and Full-Duplex-Bench (Lin et al., 2025a), commonly evaluate four key dimensions:
Timing for speaking up or interrupting: Assesses LALMs’ ability to distinguish meaningful pauses from turn-yielding moments, avoiding undesired interruptions and taking over turns appropriately.
Backchanneling: Evaluates whether LALMs backchannel at proper moments with suitable frequency, reflecting their active listening.
Turn taking: Examines whether LALMs transition smoothly between turns by recognizing boundaries, managing latency, and signaling their intent to maintain or yield the floor.
User interruption handling: Assesses LALMs’ handling of interruption, e.g., pausing and smoothly resuming the conversation.
Both use automatic evaluation metrics. Talking Turns uses supervised models trained on human dialogues (Godfrey et al., 1992) as a reference, while Full-Duplex-Bench uses metrics like response latency. However, these methods often rely on heuristics, which may be inaccurate in some cases.
5.2 Instruction Following
Instruction following is the ability to follow user-specified instructions, e.g., requirements for performing particular actions, adhering to constraints, and adjusting response styles. Effective instruction following is essential for model controllability.
LALM instruction-following evaluations typically involve three approaches: (1) adding constraints to existing LALM benchmarks not originally for instruction following, (2) synthesizing LLM instruction-following benchmarks into speech, or (3) creating new dedicated datasets. For instance, Speech-IFeval (Lu et al., 2025) introduces constraints into LALM benchmarks such as Dynamic-SUPERB Phase-2 (Huang et al., 2025a); VoiceBench (Chen et al., 2024c) synthesizes IFEval (Zhou et al., 2023a), a text-based LLM instruction-following benchmark, into speech; and URO-Bench (Yan et al., 2025) creates custom evaluation datasets.
Evaluating instruction adherence helps distinguish limitations in following instructions and deficiencies in auditory understanding or knowledge. Common evaluated constraints include length (e.g., a minimum number of words), format (e.g., responses in JSON or all caps), action (e.g., chain-of-thought reasoning (Wei et al., 2022)), style (e.g., responses in a humorous tone), and content (e.g., including a specific word). During evaluation, instruction-following rates, i.e., the frequency with which instructions are correctly followed, are measured with rule-based (Zhou et al., 2023a) or LLM-as-a-judge methods (Gu et al., 2024).
Benchmark results reveal significant gaps in LALMs compared to their LLM backbones in instruction following (Lu et al., 2025), indicating catastrophic forgetting when adapting LLMs to auditory modalities.
6 Fairness, Safety, and Trustworthiness
Despite the advancements of LALMs, their real-world deployment may pose social risks, such as perpetuating biases, generating harmful content, or spreading misinformation, if not properly evaluated and regulated. Therefore, fairness, safety, and trustworthiness must be thoroughly assessed. This section reviews works that quantify these risks to ensure the responsible and ethical use of LALMs.
6.1 Fairness and Bias
Fairness and bias are key ethical concerns for LALMs, ensuring they do not reinforce societal inequalities, discrimination, stereotypes, or biases. Such issues can be triggered by either the speech content or its non-content acoustic cues. For example, content-triggered bias may arise when LALMs translate occupation-related terms in the speech content into stereotypical gendered terms, independent of acoustic characteristics. In contrast, acoustic-triggered bias may arise when vocal cues lead the model to associate a speaker’s gender with certain occupations.
Lin et al. (2024c) quantifies LALMs’ content-triggered gender biases via four tasks: speech-to-text translation, coreference resolution, sentence continuation, and question answering. In each task, gender biases and stereotypes are measured based on the models’ responses.
Conversely, Spoken Stereoset (Lin et al., 2024b) assesses acoustic-triggered bias on speakers’ gender and age. The authors sampled sentences from NLP datasets like Stereoset (Nadeem et al., 2021) and BBQ (Parrish et al., 2022), which were then rewritten in the first-person perspective with explicit gender or age indicators (e.g., “mother”) removed to ensure bias would be triggered by speaker characteristics rather than content. The modified sentences were synthesized into speech using TTS with voices of different genders and ages. These spoken sentences served as the context, and LALMs were tasked with selecting continuations from options that were stereotypical, anti-stereotypical, or unrelated to the context.
These works highlight LALMs’ social biases, which may be inherited from their training data or LLM backbones. Additionally, since social biases are multifaceted, current benchmarks cannot include all possible societal factors, emphasizing the need for further research into both model development and benchmarks to enhance fairness.
6.2 Safety
Unlike fairness and bias, which expose societal prejudices in LALMs, safety concerns focus on preventing harmful or unsafe outputs that may negatively impact individuals or society, including user discomfort or illegal activities. Current studies typically use NLP datasets with malicious queries and convert them into speech via TTS. For example, VoiceBench (Chen et al., 2024c) and Roh et al. (2025) synthesize datasets like AdvBench (Zou et al., 2023) into spoken queries, evaluating LALMs on their ability to reject them.
During evaluation, jailbreaking techniques may be employed to test models’ resistance to adversarial inputs. These include modifying speech content by inserting fictional scenarios (Shen et al., 2024) and applying auditory manipulations such as silence (Yang et al., 2025b), noise (Yang et al., 2025b; Xiao et al., 2025), accents (Roh et al., 2025; Xiao et al., 2025), and audio edits (Xiao et al., 2025; Gupta et al., 2025). Ideally, LALMs should remain robust to adversarially modified inputs and consistently reject malicious requests.
However, evaluations show that LALMs often accept malicious spoken inputs even when they can refuse similar textual ones (Chen et al., 2024c). Moreover, LALMs show considerable safety degradation compared to their LLM backbones (Yang et al., 2025b). Several jailbreaking methods can easily bypass these models (Roh et al., 2025; Xiao et al., 2025), highlighting the need for better multimodal safety alignment.
6.3 Hallucination
Hallucination occurs when a model generates nonfactual or unsupported outputs, reducing reliability and misleading users. In LALMs, hallucinations can originate from both auditory and textual modalities. While textual hallucinations can be assessed with NLP benchmarks (Li et al., 2023; Chen et al., 2024a; Bang et al., 2025), we focus on auditory-induced hallucinations.
Kuan et al. (2024a) explores LALMs’ object hallucination, where the models falsely identify objects or events absent from the auditory input. They evaluate this via two tasks: a discriminative task where LALMs determine whether a specified object exists in the audio, and a generative task where LALMs generate captions describing the audio. These captions are then evaluated for accuracy in reflecting the actual content of the audio. Despite generating accurate captions, LALMs struggle with object identification in the discriminative task, revealing challenges in object hallucination for question-answering tasks.
Leng et al. (2024) further analyzes object hallucination using the CMM benchmark, showing that overrepresented objects or events in the training data can lead LALMs to incorrectly predict their presence, even when they are absent. Additionally, the frequent co-occurrence of objects and events during training exacerbates these hallucinations.
These works highlight hallucination challenges in LALMs and call for improved training, modeling, and data handling to enhance trustworthiness.
7 Challenges and Future Directions
7.1 Data Leakage and Contamination
Creating and curating high-quality auditory data is far more difficult than for text. Consequently, many LALM benchmarks rely on existing auditory corpora (Panayotov et al., 2015a; Kim et al., 2019; Gemmeke et al., 2017) rather than collecting new data. This raises concerns about data leakage, since models may have seen these datasets during training (Deng et al., 2024; Zhou et al., 2023b; Jacovi et al., 2023), undermining evaluation reliability. The risk grows when large-scale web-crawled data (Radford et al., 2023; He et al., 2024) are used for training without rigorous filtering.
7.2 Inclusive Evaluation Across Linguistic, Cultural, and Communication Diversity
While current benchmarks cover major languages like English and Mandarin (Huang et al., 2025a; Yan et al., 2025), many overlook crucial aspects such as low-resource languages (Magueresse et al., 2020) and code-switching (Doğruöz et al., 2021; Sitaram et al., 2019). Although these have been explored in traditional speech technologies (Khare et al., 2021; Bhogale et al., 2024; Liu et al., 2024; Yang et al., 2024b), they remain underexamined in LALMs. This limited coverage fails to capture the full linguistic diversity of human communication, as different languages possess unique characteristics (Evans and Levinson, 2009; Bickel, 2014).
Cultural factors, shaped by historical and social contexts, influence dimensions like moral norms (Graham et al., 2016; Saucier, 2018) and are essential for evaluation. As LALMs extend to diverse cultures (Yang et al., 2024a; Wang et al., 2025b), evaluation frameworks must also expand.
Along with language and culture, communication patterns also matter. While some work covers speech variations like accents, underrepresented groups such as people with speech disorders (e.g., dysarthria (Kent et al., 1999; Kim et al., 2008)) are often overlooked, as current LALMs have limited familiarity with their unique speech patterns, which affects fair and accurate understanding.
To develop fair and broadly applicable LALMs, future evaluations should carefully consider linguistic, cultural, and communicative diversity.
7.3 Safety Evaluation Unique to Auditory Modalities
Current LALM safety evaluations Section˜6.2 mainly target harmful content in model outputs, often overlooking risks inherent to auditory modalities. Auditory cues such as tone, emotion, and voice quality can also influence user experience and raise concerns if uncontrolled. For instance, even harmless content can discomfort users if spoken harshly or sarcastically, and the presence of annoying noises can also cause irritation. Thus, safety should cover auditory comfort, not just content harmlessness.
Most benchmarks focus on content toxicity but seldom assess auditory-specific safety. Addressing these issues is vital for applications like voice assistants (Pias et al., 2024; Mari et al., 2024), where vocal manner greatly affects user trust and comfort. Future work should jointly consider vocal tone, noise, and other paralinguistic factors to ensure safe, user-friendly interactions.
7.4 Unified Evaluation of Harmlessness and Helpfulness
Harmlessness and helpfulness in LALMs refer to safety and fairness, and the ability to assist users, respectively. Ideally, these two properties should be enhanced together; however, in practice, they often conflict (Bai et al., 2022). For example, a model that always refuses to answer is safe but unhelpful, as it fails to assist users. A recent study (Lin et al., 2025b) shows that post-training aimed at enhancing harmlessness can reduce helpfulness, causing models to reject queries even when no safety or privacy issues exist. This tension highlights the need for a unified evaluation framework that considers both aspects simultaneously.
Existing harmlessness benchmarks section˜6 rarely include helpfulness, limiting understanding of their trade-offs and offering limited guidance for balancing them effectively. Thus, developing a joint evaluation framework is a key future direction.
7.5 Personalization Evaluation
Personalization enables models to adapt to individual users by incorporating private information like users’ voices and preferences, supporting applications such as personalized voice assistants.
While traditional speech technologies have explored personalization (Lee et al., 2024; Joseph and Baby, 2024), it remains underdeveloped for LALMs. Unlike recent progress in LLM personalization (Tan et al., 2024, 2025; Zhang et al., 2024), LALM personalization is more complex due to the auditory dimension: LALMs must adapt to user-specific knowledge, as text LLMs do, but also become familiar with users’ voice characteristics and speaking habits, and adjust their own speaking style to match user preferences. Such complexity necessitates the development of specialized evaluations to fully assess LALM personalization, making it a valuable area for future investigation.
8 Conclusion
Holistic evaluation of LALMs is as crucial as modeling and training in advancing the field. This survey reviews existing evaluation frameworks and proposes a taxonomy categorizing current progress into four important research areas, reflecting the diverse expectations of LALM capabilities. We present a thorough overview of the literature, highlighting challenges and future directions, such as data contamination, inclusivity, auditory-specific safety, and personalization. We hope this survey provides clear guidelines for researchers and stimulates further advancements in LALM evaluation.
Limitations
We acknowledge a few limitations in this paper. First, the scope of our taxonomy is based on existing evaluation frameworks and benchmarks, meaning it does not cover all possible real-world auditory tasks. The auditory modalities are inherently complex, with a wide range of tasks and applications that cannot be exhaustively covered. As LALMs continue to evolve, new capabilities and applications will emerge, leading to growing expectations for these models. Consequently, the evaluation landscape will likely expand and shift, requiring our taxonomy to be updated and adapted to include these new tasks and applications. We will continue to follow the advancements in this field and adjust our taxonomy accordingly to reflect these developments.
Second, this survey primarily focuses on current benchmarks used to evaluate LALMs’ performance across various aspects. As a result, it does not put much emphasis on more basic or traditional evaluation methods, such as subjective assessments of speech generation quality (e.g., Mean Opinion Score), which are commonly used to evaluate model-generated audio. While these methods are valuable in certain applications, they fall outside the scope of this paper, which aims to provide a comprehensive overview of more advanced and specialized benchmarks.
Acknowledgments
We thank the reviewers for their valuable feedback, which have helped us improve the paper. To enhance the comprehensiveness of our survey, we have added further discussion in the appendix. This work was supported by the Ministry of Education (MOE) of Taiwan under the project Taiwan Centers of Excellence in Artificial Intelligence, through the NTU Artificial Intelligence Center of Research Excellence (NTU AI-CoRE)
Appendix A Detailed Categorization of the Surveyed Papers
The complete categorization of the surveyed papers, based on the proposed taxonomy section˜2, is presented in Figure 3. Please note that widely used corpora for fundamental auditory processing tasks, such as speech recognition and audio captioning, are excluded from this categorization due to the extremely large number of such resources. Including them would make the figure overly detailed and cumbersome. For reference, we provide examples of these fundamental tasks and their corresponding resources in Appendix C.
From Figure 3, it is evident that the current focus of LALM evaluations predominantly centers on auditory processing tasks section˜3.2, underscoring their importance to the community. While these tasks are valuable, they should not be seen as the sole consideration when evaluating models for real-world applications. A more diverse and comprehensive evaluation scope is crucial to ensure a fuller understanding of their potential and shortcomings.
Appendix B Brief Summary for Benchmarks Discussed in the Main Text
In this section, we summarize the key features and evaluation metrics of the benchmarks presented in Figure 2, aiming to guide researchers in selecting benchmarks suitable for their own use cases. The details are provided in Table 5 and Table 6.
| Auditory Tasks | Common Datasets |
| Audio Tasks | |
| Audio Captioning | AudioCaps Kim et al. (2019) Clotho Drossos et al. (2020) |
| Audio Classification | ESC-50 Piczak (2015b) AudioSet Gemmeke et al. (2017) |
| Vocal Sound Classification | VocalSound (Gong et al., 2022) |
| Speech Tasks | |
| Automatic Speech Recognition | LibriSpeech Panayotov et al. (2015b) AISHELL-1 Bu et al. (2017) Common Voice Ardila et al. (2020) |
| Speaker Identification | VoxCeleb2 Chung et al. (2018) CN-Celeb Fan et al. (2020) |
| Text-to-Speech | LJSpeech Ito and Johnson (2017) VCTK Veaux et al. (2017) LibriTTS Zen et al. (2019) |
| Speech Emotion Recognition | IEMOCAP Busso et al. (2008) CREMA-D Cao et al. (2014) |
| Language Identification | VoxLingua107 Valk and Alumäe (2021) FLEURS Conneau et al. (2023) |
| Speech Translation | CoVoST 2 Wang et al. (2021) MuST-C Di Gangi et al. (2019) |
| Speech Diarization | LibriMix Cosentino et al. (2020) |
| Keyword Spotting | Speech Command Warden (2018) |
| Music Tasks | |
| Music Captioning Text-to-Music | MusicCaps Agostinelli et al. (2023) Song Describer Dataset Manco et al. (2023) MidiCaps Melechovsky et al. (2024) |
| Music Transcription | MAESTRO Hawthorne et al. (2019) |
| Instrument Classification | NSynth (Engel et al., 2017) |
| Genre Classification | FMA (Defferrard et al., 2017) GTZAN (Tzanetakis and Cook, 2002) |
| Dataset | # of Benchmarks Using the Dataset |
| AudioCaps | 6 |
| Clotho | 5 |
| ESC-50 | 3 |
| AudioSet | 8 |
| VocalSound | 2 |
| LibriSpeech | 7 |
| Common Voice | 8 |
| VoxCeleb (1&2) | 3 |
| LJSpeech | 3 |
| VCTK | 4 |
| LibriTTS | 2 |
| IEMOCAP | 4 |
| CREMA-D | 2 |
| VoxLingua107 | 1 |
| CoVoST 2 | 2 |
| LibriMix | 1 |
| Speech Command | 2 |
| MusicCaps | 4 |
| Song Describer Dataset | 3 |
| MAESTRO | 1 |
| NSynth | 2 |
| FMA | 2 |
| Benchmark | Real Data | Synthetic Data |
| SALMon (Maimon et al., 2025) | ✓ | ✓ |
| EmphAssess (Seyssel et al., 2024) | ✓ | |
| Dynamic-SUPERB (Huang et al., 2024a) | ✓ | ✓ |
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a) | ✓ | ✓ |
| AIR-Bench (Yang et al., 2024c) | ✓ | ✓ |
| AudioBench (Wang et al., 2025a) | ✓ | ✓ |
| MuChoMusic (Weck et al., 2024) | ✓ | |
| ZeroSpeech 2021 (Nguyen et al., 2020) | ✓ | ✓ |
| CSZS (Huang et al., 2024b) | ✓ | |
| sStoryCloze & tStoryCloze (Mostafazadeh et al., 2017) | ✓ | |
| MMAU (Sakshi et al., 2025) | ✓ | ✓ |
| Audiopedia (Penamakuri et al., 2025) | ✓ | |
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ |
| VoxEval (Cui et al., 2025) | ✓ | |
| CompA (Ghosh et al., 2024b) | ✓ | ✓ |
| SAKURA (Yang et al., 2025a) | ✓ | |
| URO-Bench (Yan et al., 2025) | ✓ | ✓ |
| Audio Entailment (Deshmukh et al., 2025a) | ✓ | |
| Wang et al. (2025d) | ✓ | ✓ |
| StyleTalk (Lin et al., 2024a) | ✓ | |
| SD-Eval (Ao et al., 2024) | ✓ | ✓ |
| VoxDialogue (Cheng et al., 2025) | ✓ | |
| Talking Turns (Arora et al., 2025b) | ✓ | |
| Full-Duplex-Bench (Lin et al., 2025a) | ✓ | ✓ |
| Speech-IFEval (Lu et al., 2025) | ✓ | |
| Lin et al. (2024c) | ✓ | ✓ |
| Spoken Stereoset (Lin et al., 2024b) | ✓ | |
| Yang et al. (2025b) | ✓ | |
| Roh et al. (2025) | ✓ | |
| Kuan et al. (2024a) | ✓ | |
| CMM (Leng et al., 2024) | ✓ |
Appendix C Examples of General Auditory Processing Tasks and Resources
Table 1 lists representative auditory processing tasks and their associated resources. As foundational components of auditory processing, these tasks are well-suited for adaptation in LALM evaluation, as discussed in Section˜3.2.
Appendix D Overlap between Common Corpora and Existing Benchmarks
In Table 1, we list several corpora that are commonly used in the community. However, as discussed in section˜7.1, evaluation benchmarks may face risks of data leakage and contamination if they heavily rely on these existing resources. To provide a quantitative view of this issue, we report detailed statistics on the number of benchmarks discussed in the main text (Figure 2) that make use of the datasets in Table 1.
From Table 2, we observe that certain corpora are used with particularly high frequency, which reinforces concerns about data contamination, especially when web-scraped data are incorporated into training without rigorous filtering. We emphasize that this issue should be taken seriously.
Appendix E Utilization of Synthetic Auditory Data in Benchmarks in the Main Text
Table 3 summarizes whether the benchmarks in Figure 2 use real or synthetic auditory data. Synthetic audio is mainly employed to (1) generate data difficult to obtain in real settings, such as controlled stress or specific sound effects, and (2) verbalize task instructions or dialogues consistently.
Appendix F Dynamics in Full-Duplex Dialogues
In this section, we briefly introduce the dynamics discussed in section˜5.1.2. Turn-taking (Sacks et al., 1974) is a fundamental aspect of conversational organization, where speakers alternate turns to speak, ensuring only one person talks at a time. This process is complex, involving various behaviors that help facilitate smooth transitions between speakers. For example, speakers often signal the end of their turn through clear cues, allowing the listener to recognize when they are yielding the floor (Duncan, 1972; Duncan and Fiske, 2015). Furthermore, turn-taking conventions may be shaped by cultural factors (Sidnell, 2007), which influence how and when speakers take their turns due to linguistic and social differences. Understanding and modeling these behaviors are essential steps toward achieving natural and effective communication in both human-human and human-AI interactions.
Backchanneling involves the listener’s use of phatic expressions that signal active listening and attentiveness to the speaker (Fujie et al., 2005). These verbal cues, such as “yeah,” “I see,” or “uh-huh,” along with non-verbal cues like nodding, serve as feedback, showing sympathy, agreement, or understanding. By offering such responses, listeners help maintain the flow of conversation without interrupting the speaker. This behavior not only fosters a sense of connection but also enhances the speaker’s feeling of being heard and understood, contributing to a more interactive and supportive dialogue. As such, backchanneling plays a crucial role in sustaining conversation dynamics and promoting positive communicative exchanges.
Speaker overlap refers to the simultaneous speech of multiple speakers, while speaker interruption occurs when one speaker interjects during another’s turn, which breaks the turn-taking principles (Gravano and Hirschberg, 2012). These phenomena are complex: they can be competitive, reflecting hostility or dominance (West, 1979; Orcutt and Harvey, 1985), or they can be neutral or supportive, helping to maintain and coordinate the flow of dialogue (Goldberg, 1990; Jefferson, 1986; Gervits and Scheutz, 2018). Despite their varying forms, both overlap and interruption are natural components of human conversation.
Appendix G Input/Output Modalities of the Surveyed Works
Our proposed taxonomy Section˜2 is organized by the evaluation objectives of the surveyed works rather than by the modalities they cover. Nevertheless, modality information is essential for researchers seeking benchmarks suited to models specialized in particular modalities. Thus, we provide the input/output modality details in Tables 7, 8, 9, and 10, corresponding to the categories of General Auditory Awareness and Processing Section˜3, Knowledge and Reasoning Section˜4, Dialogue-oriented Ability Section˜5, and Fairness, Safety, and Trustworthiness Section˜6, respectively. These tables are compiled based on the original papers of the surveyed works.
Please note that due to unique evaluation designs, some benchmarks do not produce explicit “outputs” but instead rely on input likelihood comparisons or similarity measures with specific instances. This absence of outputs is clearly indicated in the tables.
Appendix H Comparison between End-to-end LALMs and Cascaded Systems
| Type | Benchmark List |
| Cascaded Systems Outperform | Dynamic-SUPERB (Huang et al., 2024a) |
| MMAU (Sakshi et al., 2025) | |
| VoiceBench (Chen et al., 2024c) | |
| VoxEval (Cui et al., 2025) | |
| SAKURA (Yang et al., 2025a) | |
| SD-Eval (Ao et al., 2024) | |
| Speech-IFEval (Lu et al., 2025) | |
| Yang et al. (2025b) | |
| E2E LALMs Outperform | SALMon (Maimon et al., 2025) |
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a)1 | |
| AIR-Bench (Yang et al., 2024c)2 | |
| StyleTalk (Lin et al., 2024a)3 |
1
Overall, the top-performing E2E LALMs outperform cascaded systems, except for a small number of tasks where their performance is slightly inferior.
2
The top-performing E2E LALMs significantly outperform cascaded systems on the foundation benchmark, with only a slight drop in performance on the chat benchmark.
3
The top-performing E2E LALMs significantly outperform cascaded systems across various metrics, except for one where they slightly underperform.
Table 4 presents a comparison between end-to-end (E2E) LALMs and cascaded systems that couple LLMs with modules such as speech recognition, evaluated across the benchmarks in Figure 2. Cascaded systems generally perform better on benchmarks emphasizing content- or semantics-based tasks, which require only limited integration of non-semantic auditory cues. These tasks align well with the strengths of cascaded pipelines. For instance, VoiceBench (Chen et al., 2024c) and VoxEval (Cui et al., 2025) primarily assess world knowledge and reasoning abilities adapted from text-based datasets (e.g., MMLU (Hendrycks et al., 2021)), where textual representations alone are sufficient, thus giving cascaded approaches a clear advantage.
Conversely, certain benchmarks highlight the advantages of E2E LALMs. For example, SALMon (Maimon et al., 2025) requires fine-grained auditory perception that cascaded systems struggle to replicate. This suggests the importance of future benchmarks that place greater emphasis on nuanced auditory understanding and reasoning, where LALMs hold stronger potential to excel.
Appendix I Common Metrics for LALMs
As previously noted, this paper primarily focuses on benchmarks for evaluating LALMs across diverse aspects, rather than on basic or traditional evaluation metrics. In this section, we briefly introduce some of the most commonly used metrics.
To assess the generation quality and naturalness of auditory outputs, the most widely adopted measure is the Mean Opinion Score (MOS), which relies on human annotators’ judgments to provide subjective quality assessments. While effective, MOS collection can be costly and time-consuming. To mitigate this, proxy models such as UTMOS (Saeki et al., 2022) have been proposed to automatically predict MOS scores, offering a more efficient alternative to direct human evaluation.
For evaluating the content of generated auditory outputs, several additional metrics are commonly used. In cases where models can produce both speech and text, Character Error Rate (CER) and Word Error Rate (WER) measure the consistency between generated textual outputs and transcriptions of generated speech (via an ASR system), thereby quantifying alignment across modalities.
Beyond surface-level alignment, other metrics assess the semantic quality of responses. Widely used measures include ROUGE (Lin, 2004) and BERTScore (Zhang et al., 2020), which evaluate semantic overlap between model outputs and references. More recently, LLM-as-a-judge (Gu et al., 2024) has been increasingly adopted to provide flexible, criterion-driven evaluations tailored to researchers’ specific needs.
Appendix J Information on AI Assistance
We acknowledge the assistance of GPT-4.1-mini in refining the paper and improving its clarity.
| Benchmark | Features | Metrics |
| General Auditory Awareness and Processing | ||
| SALMon (Maimon et al., 2025) | Challenging benchmark for nuanced auditory awareness of semantic, paralinguistic, and acoustic information. | Likelihood-based Comparison |
| EmphAssess (Seyssel et al., 2024) | Benchmark testing awareness of prosodic features, requiring models to preserve them during speech-to-speech translation. | Precision, Recall, F1 |
| Dynamic-SUPERB (Huang et al., 2024a) | The first benchmark covering audio, speech, and music with 55 multiple-choice QA tasks. | Accuracy |
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a) | An expanded version of Dynamic-SUPERB, currently the largest benchmark in the auditory processing category, featuring 180 tasks including open-ended ones. | Accuracy, TER, MSE, KTAU, LCC, SRCC, ERR, Miss Time, WER, Sacre BLEU, MER, IoU, F1, MEDAE, Angle Diff, Abs Diff, PER, PCC, DER, CER, POS |
| AIR-Bench (Yang et al., 2024c) | Benchmark covering audio, speech, and music with both multiple-choice and open-ended questions. | Accuracy, LLM-as-a-judge |
| AudioBench (Wang et al., 2025a) | Benchmark focusing on speech and audio, comprising 8 tasks and 26 datasets. | Word Error Rate, LLM-as-a-judge, METEOR |
| MuChoMusic (Weck et al., 2024) | In-depth investigation of music-oriented knowledge and processing abilities. | Accuracy, Instruction-following rate |
| Knowledge and Reasoning | ||
| ZeroSpeech 2021 (Nguyen et al., 2020) | Evaluating linguistic knowledge (lexical and syntactic understanding), primarily in English. | Likelihood-based Comparison, Similarity Comparison |
| CSZS (Huang et al., 2024b) | Assessing semantic and syntactic knowledge in code-switching scenarios (English-Mandarin, English-French, English-Spanish). | Likelihood-based Comparison |
| sStoryCloze & tStoryCloze (Mostafazadeh et al., 2017) | Benchmarks for semantic coherence in spoken story continuation. Focused on English. | Likelihood-based Comparison |
| MMAU (Sakshi et al., 2025) | Expert knowledge and advanced reasoning (e.g., temporal, causal reasoning across audio, speech, music). | Accuracy |
| Audiopedia (Penamakuri et al., 2025) | Knowledge-intensive QA benchmark for evaluating world knowledge of entities. | Accuracy, F1 |
| VoiceBench (Chen et al., 2024c) | Comprehensive benchmark for world knowledge, instruction following, safety, and robustness to acoustic variations. | LLM-as-a-judge, Accuracy, Refusal Rate |
| VoxEval (Cui et al., 2025) | Spoken version of MMLU considering speaker, style, and audio quality variations. | Accuracy |
| CompA (Ghosh et al., 2024b) | Compositional reasoning of temporal order and attribute binding of sound events and captions. | Self-defined Text, Audio, and Group Scores |
| SAKURA (Yang et al., 2025a) | Multi-hop reasoning benchmark integrating knowledge and auditory information (gender, emotion, language, animal sounds). | Accuracy |
| URO-Bench (Yan et al., 2025) | Benchmark of knowledge, safety, and instruction-following (English, Chinese, code-switching). | Word Error Rate, Character Error Rate, LLM-as-a-judge, UTMOS, Latency |
| Audio Entailment (Deshmukh et al., 2025a) | Deductive reasoning from auditory information. | Accuracy, Precision, Recall, F1 |
| Wang et al. (2025d) | Joint reasoning integrating acoustic and speech information. | LLM-as-a-judge |
| Benchmark | Features | Metrics |
| Dialogue-oriented Ability | ||
| StyleTalk (Lin et al., 2024a) | Earliest benchmark for affective and contextual conversational abilities of LALMs. Primarily focuses on speaker emotion and speaking styles. | BLEU, ROUGE, METEOR, BERT Score, F1, Human Evaluation (with A/B test) |
| SD-Eval (Ao et al., 2024) | Benchmark for conversational abilities of LALMs, covering speaker emotion, accent, age, and environmental sounds. | LLM-as-a-judge, ROUGE-L, BLEU, METEOR, BERT Score, Human Evaluation |
| VoxDialogue (Cheng et al., 2025) | Benchmark for conversational abilities of LALMs, further expanding the scope to 12 auditory attributes. | BLEU, ROUGE-L, METEOR, BERT Score, F1, LLM-as-a-judge |
| Talking Turns (Arora et al., 2025b) | Benchmark for evaluating the turn-taking dynamics of LALMs in full-duplex dialogues. | Automatic judgment with an internally trained judge model |
| Full-Duplex-Bench (Lin et al., 2025a) | Benchmark for pause handling, backchanneling, turn-taking, and interruption management in full-duplex dialogues. | Takeover Rate, Backchannel Frequency, Jensen-Shannon Divergence, Latency, LLM-as-a-judge |
| Speech-IFEval (Lu et al., 2025) | Benchmark specifically tailored for instruction following, disentangling instruction-following from speech perception. Can also analyze the degree of catastrophic forgetting in LALMs compared with their LLM backbones. | Word Error Rate, Accuracy, LLM-as-a-judge, Instruction-following Rate |
| Fairness, Safety, and Trustworthiness | ||
| Lin et al. (2024c) | Quantifying the gender bias of LALMs through four tasks. | Accuracy, F1, F1 Differences, Language Modeling Score, Stereotypical Score, Idealized Context Association Tests Score, Instruction Following Rate, Bias Score |
| Spoken Stereoset (Lin et al., 2024b) | Evaluating social bias in LALMs, including gender and age. | Speech Language Instruction Following Score, Speech Language Modeling Score, Speech Language Bias Score |
| Yang et al. (2025b) | Quantifying safety issues of LALMs under jailbreaking attempts. | Attack Success Rate by Attempt, Attack Success Rate by Questions |
| Roh et al. (2025) | Evaluating safety alignment of LALMs under adversarial multilingual and multi-accent audio jailbreaks. | Jailbreak Success Rate, Word Error Rate, Accuracy |
| Kuan et al. (2024a) | Benchmark for object hallucination in LALMs, covering both discriminative and generative tasks. | Accuracy, Precision, Recall, F1, Word Error Rate, Self-defined Metrics |
| CMM (Leng et al., 2024) | Benchmark for quantifying object hallucination and identifying its potential causes. Includes vision modality as well. | Perception Accuracy, Hallucination Resistance |
| General Auditory Awareness and Processing | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| SALMon Maimon et al. (2025) | ✓ | ✓ | Likelihood-based evaluation. No output modality. | |||||
| Wu et al. (2024b) | ✓ | ✓ | ✓ | |||||
| EmphAssess (Seyssel et al., 2024) | ✓ | ✓ | ||||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||
| Dynamic-SUPERB (Huang et al., 2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Dynamic-SUPERB Phase-2 (Huang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| AIR-Bench (Yang et al., 2024c) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| AudioBench (Wang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||
| FinAudio (Cao et al., 2025) | ✓ | ✓ | ✓ | |||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| MAE (Chen et al., 2024b) | ✓ | ✓ | ✓ | ✓ | ||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||
| Wang et al. (2025b) | ✓ | ✓ | ✓ | |||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Audio-FLAN (Xue et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
| QualiSpeech (Wang et al., 2025c) | ✓ | ✓ | ✓ | |||||
| EvalSIFT (Pandey et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||
| ASR-EC (Wei et al., 2024) | ✓ | ✓ | ✓ | |||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||
| BEANS-Zero (Robinson et al., 2025) | ✓ | ✓ | ✓ | |||||
| Knowledge and Reasoning | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| ZeroSpeech 2021 (Nguyen et al., 2020) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| CSZS (Huang et al., 2024b) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| sStoryCloze (Hassid et al., 2023) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| tStoryCloze (Hassid et al., 2023) | ✓ | Likelihood-based evaluation. No output modality. | ||||||
| BabySLM (Lavechin et al., 2023) | ✓ | ✓ | Likelihood-based evaluation. No output modality. | |||||
| CompA (Ghosh et al., 2024b) | ✓ | ✓ | Similarity-based evaluation on text and audio inputs. | |||||
| MMAU (Sakshi et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| Audiopedia (Penamakuri et al., 2025) | ✓ | ✓ | ✓ | |||||
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ | ✓ | |||||
| VoxEval (Cui et al., 2025) | ✓ | ✓ | ||||||
| SAKURA (Yang et al., 2025a) | ✓ | ✓ | ✓ | ✓ | ||||
| URO-Bench (Yan et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Audio Entailment (Deshmukh et al., 2025a) | ✓ | ✓ | ✓ | |||||
| ADU-Bench (Gao et al., 2024) | ✓ | ✓ | ✓ | |||||
| SAGI (Bu et al., 2024) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| MuChoMusic (Weck et al., 2024) | ✓ | ✓ | ✓ | |||||
| RUListening (Zang et al., 2025) | ✓ | ✓ | ✓ | |||||
| OpenMU-Bench (Zhao et al., 2024) | ✓ | ✓ | ✓ | |||||
| Gong et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ✓ | |||
| CompA-R (Ghosh et al., 2024a) | ✓ | ✓ | ✓ | |||||
| OpenAQA (Gong et al., 2024b) | ✓ | ✓ | ✓ | |||||
| Clotho-AQA (Lipping et al., 2022) | ✓ | ✓ | ✓ | |||||
| SLU-GLUE (Li et al., 2024) | ✓ | ✓ | ✓ | |||||
| SpeechCaps (Huang et al., 2025b) | ✓ | ✓ | ✓ | |||||
| Wang et al. (2025d) | ✓ | ✓ | ✓ | ✓ | ||||
| Deshmukh et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||
| Dialogue-oriented Ability | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| StyleTalk (Lin et al., 2024a) | ✓ | ✓ | ✓ | |||||
| SD-Eval (Ao et al., 2024) | ✓ | ✓ | ✓ | ✓ | ||||
| VoxDialogue (Cheng et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Talking Turns (Arora et al., 2025b) | ✓ | ✓ | ||||||
| Full-Duplex-Bench (Lin et al., 2025a) | ✓ | ✓ | ||||||
| Li et al. (2025) | ✓ | ✓ | ||||||
| ContextDialog (Kim et al., 2025) | ✓ | ✓ | ✓ | |||||
| ADU-Bench (Gao et al., 2024) | ✓ | ✓ | ✓ | |||||
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ | ✓ | |||||
| URO-Bench (Yan et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Speech-IFeval (Lu et al., 2025) | ✓ | ✓ | ✓ | |||||
| S2S-Arena (Jiang et al., 2025) | ✓ | ✓ | ||||||
| EvalSIFT (Pandey et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Fairness, Safety, and Trustworthiness | ||||||||
| Benchmark | Input Modalities | Output Modalities | ||||||
| Text | Audio | Speech | Music | Text | Audio | Speech | Music | |
| Lin et al. (2024c) | ✓ | ✓ | ✓ | |||||
| Spoken Stereoset (Lin et al., 2024b) | ✓ | ✓ | ✓ | |||||
| VoiceBench (Chen et al., 2024c) | ✓ | ✓ | ✓ | |||||
| Yang et al. (2025b) | ✓ | ✓ | ✓ | ✓ | ||||
| Roh et al. (2025) | ✓ | ✓ | ✓ | |||||
| AdvBench-Audio (Kang et al., 2025) | ✓ | ✓ | ✓ | |||||
| Xiao et al. (2025) | ✓ | ✓ | ✓ | |||||
| Gupta et al. (2025) | ✓ | ✓ | ✓ | |||||
| Hughes et al. (2024) | ✓ | ✓ | ✓ | |||||
| URO-Bench (Yan et al., 2025) | ✓ | ✓ | ✓ | ✓ | ||||
| Kuan et al. (2024a) | ✓ | ✓ | ✓ | ✓ | ||||
| CMM (Leng et al., 2024) | ✓ | ✓ | ✓ | |||||
| Kuan and Lee (2025) | ✓ | ✓ | ✓ | |||||